OpenStudy (anonymous):

Jenni_Fowler question: (a+2/a^2+a)/(a+2/a^3) simplify show work this

OpenStudy (anonymous):

\[\frac{(a+2)}{a^2+a} \div \frac{(a+2)}{a^3}\]

OpenStudy (anonymous):

dividing is same as multiplying by reciprocal

OpenStudy (anonymous):

\[\frac{a+2}{a^2+a} \times \frac{a^3}{a+2}\]

OpenStudy (anonymous):

a+2 cancels \[\frac{a^3}{a^2+a}\]

OpenStudy (anonymous):

can be simplified further. take out a factor of a to get a^2/ (a+1)